# Refuse Whenever You Feel Unsafe: Improving Safety in LLMs via Decoupled Refusal Training

### v1

@@ Abstract
This study addresses a critical gap in safety tuning practices for Large Language Models (LLMs) by identifying and tackling a refusal position bias within safety tuning data, which compromises the models’ ability to appropriately refuse generating unsafe content. We introduce a novel approach, Decoupled Refusal Training (DeRTa), designed to empower LLMs to refuse compliance to harmful prompts at any response position, significantly enhancing their safety capabilities. DeRTa incorporates two novel components: (1) Maximum Likelihood Estimation (MLE) with Harmful Response Prefix, which trains models to recognize and avoid unsafe content by appending a segment of harmful response to the beginning of a safe response, and (2) Reinforced Transition Optimization (RTO), which equips models with the ability to transition from potential harm to safety refusal consistently throughout the harmful response sequence. Our empirical evaluation, conducted using LLaMA3 and Mistral model families across six attack scenarios, demonstrates that our method not only improves model safety without compromising performance but also surpasses well-known models such as GPT-4 in defending against attacks. Importantly, our approach successfully defends recent advanced attack methods (e.g., CodeAttack Ren et al. (2024)) that have jailbroken GPT-4 and LLaMA3-70B-Instruct.  [^1] Our code and data can be found at https://github.com/RobustNLP/DeRTa
WARNING: This paper contains unsafe model responses.
(a) Vanilla LLMs
(b) Our LLMs
Figure 1: Vanilla LLMs with standard safety tuning tend to continue generating unsafe responses once they start doing so. In contrast, LLMs enhanced with our approach can recognize and halt the generation of unsafe content when they detect potential risks.
@@ 1 Introduction
Large Language Models (LLMs) exhibit a level of intelligence that is both impressive and ever-evolving (OpenAI, 2023; Anthropic, 2024; Meta, 2024). However, this remarkable capacity also acts as a double-edged sword, underscoring the importance of ensuring their safety. To address this, researchers have implemented various strategies to align LLMs with human ethics (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022b; Touvron et al., 2023). Despite these efforts, the challenge of rendering LLMs completely safe remains, as new safety risks continually emerge. These include jailbreak attacks (Wei et al., 2024), deceptive alignment (Hubinger et al., 2024), jailbreak fine-tuning (Qi et al., 2024b; Yang et al., 2023; Halawi et al., 2024), and adversarial attacks (Zou et al., 2023b). Notably, jailbreak attacks have garnered significant attention due to their ability to circumvent protections with simple prompts, eliminating the need for any tuning or insider knowledge.
Recent research has extensively focused on addressing jailbreak attacks through various strategies, such as prompt-based defense (Xie et al., 2023), input perturbation (Robey et al., 2023), jailbreak detection (Inan et al., 2023), priority training (Wallace et al., 2024), and representation engineering (Zou et al., 2023a). Despite these advancements in methodologies and models to improve model safety, the influence of safety tuning data remains inadequately explored.
To bridge the gap, we identify a refusal position bias in the safety tuning data, which hampers the ability of the tuned LLMs to learn how to refuse effectively. Making a refusal decision before generating the response content leads to two significant shortcomings: (1) there is a lack of necessary information for making a refusal decision, and (2) there is no mechanism to incorporate refusal at later stages of the response. Based on these observations, we propose a novel safety tuning method called Decoupled Refusal Training (DeRTa), to explicitly train LLMs to refuse compliance at any response position by embedding the constructed harmful responses. Concretely, our approach introduces two novel components:
@@
We evaluate our approach using two prominent model families: LLaMA3 (8B and 70B) (Meta, 2024) and Mistral (7B and 8$\times$7B) (Jiang et al., 2023) across six attack scenarios. Experimental results show that our method not only boosts model safety without sacrificing performance but also surpasses notable models including GPT-4 and the instructional variants of LLaMA3-70B in attack defending. Both quantitative and qualitative assessments support our assertion that our strategy effectively arms LLMs with the ability to recognize and halt the generation of unsafe content when they detect potential risks.
@@ 2 Methodology
In this section, we identify an important issue associated with the safety data – a refusal position bias that compromises the tuned models’ ability to appropriately refuse generating unsafe content. Based on the observation, we propose a novel method to enhance safety by mitigating the issue of refusal position bias.
(a) Standard Safety Tuning
(b) Ours
Figure 2: Overview of (a) the standard safety tuning and (b) ours. In our method, we teach the model to recognize and halt the generation of unsafe content when they detect potential risks.
@@ 2.1 Standard Safety Tuning
Standard safety tuning methods aim to instruct the model to generate safe responses to harmful queries (Bianchi et al., 2024; Touvron et al., 2023). Formally, given a harmful query $q$ and a safe response $r$:
$\mathcal{L}_{\text{safe}}(\theta)$ $=-\mathbb{E}_{(q,r)\sim\mathcal{D}}\log P_{\theta}(r|q)=-\mathbb{E}_{(q,r)\sim\mathcal{D}}\sum\nolimits_{i=1}^{n}\log P_{\theta}(r_{i}|q,r_{<i})$ (1)
@@
Model Ratio
LLaMA3-8B-Instruct 99.5%
LLaMA3-70B-Instruct 99.5%
Table 1: Ratio of refusal responses where the refusal tokens occur in the first 5 tokens.
@@
As shown in Figure 2(a), in the safety data, the refusal tokens such as “Sorry,” “I cannot,” and “I apologize,” consistently occur within the first few tokens of a safe response. Accordingly, LLMs tuned on these safety data tend to generate refusal tokens at the beginning of a response. The results on the SOTA open-sourced LLMs with safety tuning in Table 1 confirm our claim. The refusal positional bias may lead to the following weaknesses:
@@
    Lack of Necessary Information for Refuse Decision: The tuned model needs to make a refuse decision at the beginning of a response based on the query only, which may contain insufficient information for the decision.
@@
    Lack of a Mechanism to Refuse in Later Positions: The positional bias may lead the model to rely heavily on position-specific features. Accordingly, the tuned model tends to continue generating unsafe responses once they start doing so, compromising safety in subsequent positions.
@@ 2.2 Our Approach
To address the issues identified earlier, we have developed a method where LLMs are explicitly trained to refuse compliance at any response juncture by embedding the constructed harmful responses within the training process. As depicted in Figure 2(b), our strategy is comprised of two key components, each designed to counteract the two main concerns discussed.
@@
Initially, we incorporate a segment of the harmful response, varying in length, before the safe response. This approach provides several advantages:
@@
    Incorporating a harmful prefix enriches the query with additional context, enhancing the model’s ability to discern and avert potential threats. Despite the harmful prefix not being present during practical inference scenarios, we posit that this strategy facilitates a more robust understanding and recognition of unsafe content, thereby improving the model’s reliability and safety. The ablation study in Section 3.3 confirms our claim.
@@
    It trains the model to seamlessly transition from recognizing a potentially harmful initiation to generating a safe, appropriate response. This not only fortifies the LLM against inadvertently propagating harmful content but also equips it with the capability to navigate away from precarious contexts, ensuring the generation of benign, constructive outputs.
@@
(a) MLE with Harmful Prefix
(b) Reinforced Transition Optimization (RTO)
Figure 3: Illustrations of (a) MLE with Harmful Prefix, and (b) our RTO. The transition from harmful response to safety refusal only occurs once in MLE with Harmful Prefix (the dashed square), while in RTO we simulate the transition at every position within the full harmful response sequence.
@@
Figure 3(b) illustrates the training objectives, demonstrating a departure from the previously mentioned MLE with harmful prefix (Figure 3(a)). Instead, we simulate the transition from a harmful response to a safe refusal at every position within the entire response sequence. Consequently, LLMs trained with RTO learn the transitions $L$ times ($L$ represents the length of the harmful response) more frequently than those trained with MLE with harmful prefix. This significantly enhances their ability to proactively recognize and stop the generation of unsafe content upon detecting potential risks.
@@
where $\hat{r}_{<k}$ is the first $k$ (a random number sampled from 0 to $|\hat{r}|$) tokens of the harmful response $\hat{r}$, and “sorry” is the refusal token.
@@ 3.1 Setup
We utilize 60K uncensored samples from Evol-Instruct (Xu et al., 2023) as the SFT data for helpfulness. We use harmful instructions from BeaverTails (Ji et al., 2023) as the safety data. To build safety tuning data for our approach, we sample 3,000 instructions and obtain safe responses from GPT-3.5-turbo and harmful responses from our maliciously tuned LLaMA3-8B-Instruct. Since each instance is a triple that consists of two (query, response) pairs (i.e., (harmful query, safe response) and (harmful query, harmful response)), we complement the safety dataset to 6,000 instances for the vanilla safety tuning for fair comparison.
@@
In our experiments, we consider two representative open-source model families: LLaMA3 (8B and 70B) and Mistral (7B and 8$\times$7B). To eliminate the effect of other instruction tuning data, we conduct main experiments on the officially released raw models without instruction tuning. We set the temperature to 0 for all models, and remain the other hyperparameters as default. For tuning the models, we set the total batch size to 128, and the number of epochs to 2.
@@
We conduct a random sampling of 100 harmful questions from the Do-Not-Answer dataset (Wang et al., 2024) and another 100 from HarmBench (Mazeika et al., 2024), resulting in a total of 200 harmful questions. Our evaluation encompasses several prominent black-box attack methods, including CodeAttack (Ren et al., 2024), PAIR (Chao et al., 2023), JailbreakChat (Walkerspider, 2022), and SelfCipher (Yuan et al., 2024b). For white-box attacks, we extend our analysis beyond AutoDAN (Liu et al., 2024a) by introducing an innovative method called CompletingAttack. This approach eliminates all formatting tokens (e.g., [INST]) to render the query in a declarative format, enabling the model to complete the text. CompletingAttack achieves high success rates across all tested LLMs, such as LLaMA3-70B-instruct.
We determine the Attack Success Rate (ASR) by manually evaluating the responses generated by the target LLMs for each attack method. The ASR indicates the proportion of harmful responses generated. For this metric, we used 50 harmful queries for PAIR and AutoDAN due to their computational complexity and the full set of 200 queries for the other attack methods.
@@
We also assess the helpfulness of the targeted LLMs to determine if our approach increases safety at the expense of reducing helpfulness. To do this, we randomly select 500 examples from three sources: GSM8K (math reasoning) (Cobbe et al., 2021), MMLU (knowledge tests) (Hendrycks et al., 2021), and AlpacaEval (Li et al., 2023) (general capability). We follow the common practice to evaluate the results on AlpacaEval with GPT-4, and manually evaluate the results for the other two tasks.
More details about the experimental setup can be found in Appendix (A - C).

### v2

@@ Abstract
This study addresses a critical gap in safety tuning practices for Large Language Models (LLMs) by identifying and tackling a refusal position bias within safety tuning data, which compromises the models’ ability to appropriately refuse generating unsafe content. We introduce a novel approach, Decoupled Refusal Training (DeRTa), designed to empower LLMs to refuse compliance to harmful prompts at any response position, significantly enhancing their safety capabilities. DeRTa incorporates two novel components: (1) Maximum Likelihood Estimation (MLE) with Harmful Response Prefix, which trains models to recognize and avoid unsafe content by appending a segment of harmful response to the beginning of a safe response, and (2) Reinforced Transition Optimization (RTO), which equips models with the ability to transition from potential harm to safety refusal consistently throughout the harmful response sequence. Our empirical evaluation, conducted using LLaMA3 and Mistral model families across six attack scenarios, demonstrates that our method not only improves model safety without compromising performance but also surpasses baseline methods in defending against attacks. WARNING: This paper contains unsafe model responses.  [^1] Our code, data, and results can be found at https://github.com/RobustNLP/DeRTa.
Refuse Whenever You Feel Unsafe: Improving Safety in LLMs via Decoupled Refusal Training
Youliang Yuan1,2,4thanks: Work was done when Youliang Yuan, Wenxuan Wang, and Jen-tse Huang were interning at Tencent AI Lab.  Wenxiang Jiao2  Wenxuan Wang2,3  Jen-tse Huang2,3 Jiahao Xu2         Tian Liang2      Pinjia He1,4thanks: Pinjia He is the corresponding author.        Zhaopeng Tu2 1School of Data Science, The Chinese University of Hong Kong, Shenzhen, China 2Tencent AI Lab    3The Chinese University of Hong Kong 4Shenzhen Research Institute of Big Data, China 1youliangyuan@link.cuhk.edu.cn,hepinjia@cuhk.edu.cn 2{joelwxjiao,jwxwang,jentsehuang,jettexu,ttianliang,zptu}@tencent.com
@@ 1 Introduction
(a) Standard Safety Tuning
(b) Ours
(c) MLE with Harmful Prefix
(d) Reinforced Transition Optimization (RTO)
Figure 1: Illustration of (a) the standard safety tuning, (b) ours method, (c) MLE with Harmful Prefix, and (d) RTO. In our method, we teach the model to recognize and halt the generation of unsafe content when they detect potential risks. The transition from harmful response to safety refusal only occurs once in MLE with Harmful Prefix (the dashed square), while in RTO we simulate the transition at every position within the full harmful response sequence.
Large Language Models (LLMs) exhibit a level of intelligence that is both impressive and ever-evolving (OpenAI, 2023; Anthropic, 2024; Meta, 2024). However, this remarkable capacity also acts as a double-edged sword, underscoring the importance of ensuring their safety. To address this, researchers have implemented various strategies to align LLMs with human ethics (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022b). Despite these efforts, the challenge of rendering LLMs completely safe remains, as new safety risks continually emerge  (Zou et al., 2023b; Wei et al., 2024; Qi et al., 2024; Yang et al., 2023; Halawi et al., 2024). Notably, jailbreak attacks have garnered significant attention due to their ability to circumvent protections with simple prompts, eliminating the need for any tuning or insider knowledge.
Recent research has extensively focused on addressing jailbreak attacks through various strategies, such as prompt-based defense (Xie et al., 2023), input perturbation (Robey et al., 2023), safety decoding (Xu et al., 2024c), jailbreak detection (Inan et al., 2023), knowledge editing (Wang et al., 2024a), representation engineering (Zou et al., 2023a), latent adversarial training (Sheshadri et al., 2024), and priority training (Wallace et al., 2024). Despite these advancements in methodologies to improve model safety, the influence of safety tuning data remains inadequately explored.
Figure 2: LLMs using our approach can refuse to answer whenever they feel it is unsafe, even if they are already at a later position in the response.
To bridge the gap, we identify a refusal position bias in the safety tuning data, which hampers the ability of the tuned LLMs to learn how to refuse effectively. Making a refusal decision before generating the response content leads to two significant shortcomings: (1) there is a lack of necessary information for making a refusal decision, and (2) there is no mechanism to incorporate refusal at later stages of the response. Based on these observations, we propose a novel safety tuning method called Decoupled Refusal Training (DeRTa) (see Figure 1), to explicitly train LLMs to refuse compliance at any response position by embedding the constructed harmful responses. Concretely, our approach introduces two novel components:
@@
We evaluate our approach using two prominent model families: LLaMA3 (8B and 70B) (Meta, 2024) and Mistral (7B-v0.1 and 8$\times$7B) (Jiang et al., 2023) across six attack scenarios. Experimental results show that our method not only improves model safety without sacrificing helpfulness but also surpasses notable models including GPT-4, LLaMA3-Instruct, and all five baseline methods in attack defending. Both quantitative and qualitative assessments support our assertion that our strategy effectively arms LLMs with the ability to refuse whenever they detect potential risks.
2 Related Work
Jailbreak Attack on LLMs.
Ensuring that LLMs align with human ethics and preferences is essential to their responsible deployment (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022a; Rafailov et al., 2024). While aligning LLMs with safety data is beneficial, these models remain vulnerable to jailbreak inputs (Shen et al., 2024). Researchers have discovered that safety mechanisms can be circumvented by transforming the malicious query into semantically equivalent forms, such as ciphers (Yuan et al., 2024a), low-resource languages (Wang et al., 2024b; Deng et al., 2024; Yong et al., 2023), or code (Ren et al., 2024). Another effective jailbreak method is to frame the malicious question in a hypothesis scenario that makes it appear harmless (Chao et al., 2023; Liu et al., 2024; Wu et al., 2025). Given the high intelligence of LLMs, insights from social science (Zeng et al., 2024) and psychology (Zhang et al., 2024a) have also been applied to uncover safety issues. Moreover, techniques like adversarial suffix optimization (Zou et al., 2023b), few/many-shot attacks (Wei et al., 2023; Anil et al., 2024), multi-turn jailbreak (Li et al., 2024). According to Wei et al. (2024), the success of these attacks can be attributed to “competing objectives” and “mismatched generalization”.
Jailbreak Defense.
Current defense strategies against jailbreak attacks primarily involve safety prompts (Xie et al., 2023; Zheng et al., 2024), input perturbation (Robey et al., 2023; Cao et al., 2024), safety decoding (Xu et al., 2024c), jailbreak detection (Inan et al., 2023), representation engineering (Zou et al., 2023a; Wang et al., 2024a; Zou et al., 2024), adversarial training (Mazeika et al., 2024; Sheshadri et al., 2024), and priority training (Wallace et al., 2024). Jailbreak detection typically utilizes LLMs to identify attempted attacks (Phute et al., 2024; Zhang et al., 2024c), or involves training specialized classifiers to detect jailbreaks (Inan et al., 2023; Yuan et al., 2024b; Jain et al., 2023; Alon and Kamfonas, 2023; Hu et al., 2024; Zhang et al., 2025). Priority training methods (Zhang et al., 2024b; Lu et al., 2024) involve using strategically designed data to train LLMs to prioritize higher-ranked instructions, allowing developers to set safety prompts to the highest priority post-deployment to prevent jailbreak attempts.
In this study, we establish a connection between these vulnerabilities and a bias towards refusal positions in the tuning data, which is used to align with safety protocols. Concurrently, related work by  Qi et al. (2025); Xu et al. (2024b) has also highlighted a tendency in safety alignment to take shortcuts, specifically, alignment often prioritizes adaptations in the model’s over only its very first few output tokens. In addressing this issue, they suggest a straightforward data augmentation strategy aimed at deepening safety alignment by training with data that begins with harmful responses but eventually shifts towards safety refusals. Our research primarily diverges in two aspects: (1) we explore vulnerabilities through the lens of refusal position bias, as opposed to focusing on the generative distribution; and (2) we show that merely starting with harmful response prefixes is inadequate for countering various forms of attacks, including sophisticated methods like CodeAttack and CompletingAttack (see Figure 3 and Table 3).
@@ 3 Methodology
In this section, we identify an important issue associated with the safety data – a refusal position bias that compromises the tuned models’ ability to refuse generating unsafe content. Based on the observation, we propose a novel method to enhance safety by mitigating the refusal position bias.
@@ 3.1 Standard Safety Tuning
Standard safety tuning aims to instruct the model to generate safe responses to harmful queries (Bianchi et al., 2024; Touvron et al., 2023). Formally, given a harmful query $q$ and a safe response $r$:
$\mathcal{L}_{\text{safe}}(\theta)$ $=-\mathbb{E}_{(q,r)\sim\mathcal{D}}\log P_{\theta}(r|q)$ (1)
$=-\mathbb{E}_{(q,r)\sim\mathcal{D}}\sum\nolimits_{i=1}^{n}\log P_{\theta}(r_{i}|q,r_{<i})$
@@
Refusal Token Number Position
      (|Total Query|=800) $\leq 5^{th}$ $>5^{th}$
LLaMA3-8B-Instruct 478 2
LLaMA3-70B-Instruct 441 2
Table 1: The number of responses where refusal tokens appear within the first 5 tokens and after the first 5 tokens across six attack tasks. A small number of later refusals suggests that if the model does not refuse at the start, its safeguards can be easily bypassed.
@@
As shown in Figure 1(a), in the safety data, the refusal tokens such as “Sorry,” “I cannot,” and “I apologize,” consistently occur within the first few tokens of a safe response. Accordingly, LLMs tuned on these safety data struggle to generate refusal tokens in the later parts of a response. The results in Table 1 (and Figure 4) confirm our claim. The refusal positional bias may lead to the following weaknesses:
@@
    Lack of Necessary Information for Refuse Decision: The model needs to make a refuse decision at the beginning of a response based on the query only, which may contain insufficient information for the decision. This situation is demonstrated in the CodeAttack example shown in Figure 2.
@@
    Lack of a Mechanism to Refuse in Later Positions: The positional bias may lead the model to rely heavily on position-specific features. Accordingly, the model tends to continue generating unsafe responses once they start doing so, compromising safety in subsequent positions.
@@ 3.2 Our Approach
To address the issues identified, we have developed a method where LLMs are explicitly trained to refuse compliance at any response juncture by embedding the constructed harmful responses within the training process. As depicted in Figure 1(b), our strategy is comprised of two key components:
@@
 [^2] The harmful prefix are excluded from the loss function, so the model is not encouraged to learn patterns of “intentionally generating harmful content first, followed by safe content.”
We incorporate a segment of the harmful response, varying in length, before the safe response. This approach provides several advantages:
@@
    Incorporating a harmful prefix enriches the query with additional context, enhancing the model’s ability to discern and avert potential threats. Despite the harmful prefix not being present during practical inference scenarios, we posit that this strategy facilitates a more robust understanding of unsafe content, thereby improving the model’s safety. The ablation study in Section 4.3 confirms our claim.
@@
    It trains the model to seamlessly transition from recognizing a potentially harmful initiation to generating a safe, appropriate response. This equips the model with the capability to navigate away from precarious contexts, ensuring the generation of benign, constructive outputs.
@@
Figure 1(d) illustrates the training objectives, demonstrating a departure from the previously mentioned MLE with harmful prefix (Figure 1(c)). Instead, we simulate the transition from a harmful response to a safe refusal at every position within the entire response sequence. Consequently, LLMs trained with RTO learn the transitions $L$ times ($L$ represents the length of the harmful response) more frequently than those trained with MLE with harmful prefix. This significantly enhances their ability to proactively recognize and stop the generation of unsafe content upon detecting potential risks.
@@
[…]
where $\hat{r}_{<k}$ is the first $k$ (a random number sampled from 0 to $|\hat{r}|$) tokens of the harmful response $\hat{r}$, and “sorry” is the refusal token. Moreover, as shown in the loss, harmful tokens do not receive gradient backpropagation, which prevents the model from intentionally generating harmful content.
@@ 4.1 Setup
We utilize 60K uncensored samples from Evol-Instruct (Xu et al., 2024a) as the SFT data for helpfulness. We use harmful instructions from BeaverTails (Ji et al., 2023) as the safety data. To build safety tuning data for our approach, we sample 3,000 instructions and obtain safe responses from GPT-3.5-turbo and harmful responses from our maliciously tuned LLaMA3-8B-Instruct.
@@
We consider two representative open-source model families: LLaMA3 (8B and 70B) and Mistral (7B-v0.1 and 8$\times$7B). For large-scale models, we apply the LoRA method (Hu et al., 2022). To eliminate the effect of other instruction tuning data, we conduct main experiments on the officially released raw models without instruction tuning. For tuning the models, we set the total batch size to 128, and the number of epochs to 2.
Baselines
In our experiments, we compare our approach to several commonly used methods: vanilla safety training (Bianchi et al., 2024), GoalPriority (Zhang et al., 2024b), SoFA (Lu et al., 2024), and RecAug (Qi et al., 2025). Both our method and these baselines focus on improving safety through adjustments to the training data, without modifying the standard fine-tuning and decoding framework. Additionally, similar to our method, these approaches do not introduce any extra costs during training or inference, nor do they require the use of additional safety detectors. To further explore the impact of harmful responses within the training data, we include DPO (Rafailov et al., 2024) as another baseline for comparison.
@@
We collected 100 harmful questions each from the Do-Not-Answer dataset (Wang et al., 2024c) and HarmBench (Mazeika et al., 2024), resulting in a fixed evaluation set of 200 harmful questions. Our evaluation encompasses several prominent black-box attack methods, including CodeAttack (Ren et al., 2024), PAIR (Chao et al., 2023), JailbreakChat (Walkerspider, 2022), and SelfCipher (Yuan et al., 2024a). For white-box attacks, we extend our analysis beyond GCG (Zou et al., 2023b) [^3] Due to the computational cost limitation, we only include the results of GCG for small-scale models. and AutoDAN (Liu et al., 2024) by introducing a method called CompletingAttack. This approach eliminates all formatting tokens (e.g., [INST]) to render the query in a declarative format, enabling the model to complete the text. CompletingAttack achieves high success rates across all tested LLMs.
We determine the Attack Success Rate (ASR) by manually evaluating the responses generated by the target LLMs for each attack method, based on the evaluation criteria outlined in Appendix C. The ASR indicates the proportion of harmful responses generated. For this metric, we used a fixed subset of 50 harmful queries for PAIR and AutoDAN due to their computational complexity and the full set of 200 queries for the other attack methods.
@@
We also assess the helpfulness of the targeted LLMs to determine if our approach increases safety at the expense of reducing helpfulness. To do this, we select 500 examples from three sources: GSM8K (math reasoning) (Cobbe et al., 2021), MMLU (knowledge tests) (Hendrycks et al., 2021), and AlpacaEval (Li et al., 2023) (general capability). We follow the common practice to evaluate the results on AlpacaEval with GPT-4, and manually evaluate the results for the other two tasks.
In all evaluation experiments, we apply greedy decoding. More details about the experimental setup can be found in Appendix (A - C).